### v1

# Quantitative agreement between experiment and theory for Vibrational Circular Dichroism enhanced by electronically excited states

Abstract
Intensity enhancement in vibrational circular dichroism (VCD) arises in open-shell transition metal complexes from coupling between low-lying excited states (LLESs) and ground-state vibrational modes. In this work we apply Nafie’s vibronic coupling theory to M(II)-(-)-sparteine-Cl2 to investigate these enhancement effects. We show that the VCD intensity is extremely sensitive to the excitation energies that neither time-dependent density functional theory (TDDFT) nor state-averaged complete active space self consistent field (SA-CASSCF) calculations can predict with sufficient accuracy. We argue that instead of using more accurate quantum chemistry methods these excitation energies can be treated as parameters and optimized against experimental spectra. With this approach we obtain simulated VCD similarity scores above 0.4, a threshold considered reliable for absolute configuration assignment.
1 Introduction
Natural products and other molecules of interest for drug discovery often contain one or more chiral centers that are crucial for their biological activity. This characteristic makes it essential to be able to determine their absolute configuration (AC) or enantiomeric composition 1, 2, 3, 4, 5. In addition to electronic circular dichroism (ECD) spectroscopy, vibrational circular dichroism (VCD) spectroscopy offers a valuable complementary approach for structure determination6. VCD spectra are feature-rich, provide detailed information of the chemical bonds close to the chiral centers, and give insight into intermolecular interactions 7. Many reviews cover both experimental and theoretical VCD methods 3, 8, 2, 4, 9, 10, however, they typically focus on closed-shell systems whose electronic ground state is well separated from excited states. In those cases the VCD spectrum can be efficiently modelled using the magnetic field perturbation (MFP) theory of Stephens11 although the uncertainty in conformer energies can affect the reliability of the stereochemical assignment 12. For molecules with low-lying excited states, as is typically the case in transition metal complexes, the situation is different as the vibrational and electronic degrees of freedom of the molecule are not well-separated. Such chiral molecules are relevant in chiral material science and asymmetric catalysis 13, 14 and there is therefore a growing interest in VCD spectroscopy of molecules containing transition metals 15, 16, 17, 18, 19 that show notably intense VCD peaks.
The phenomenon of enhanced VCD signals in transition metal complexes has been known since 1980, when the first experimental VCD spectra for (--)-sparteine complexes M(II)(sp)Cl2 (M = Zn, Co, Ni) were reported in a narrow region of CH stretching modes 20. Later He et al. extended the measurement to the fingerprint region, where the open-shell Co(II) and Ni(II) have about ten times higher VCD intensity compared to the closed-shell Zn complex 21. Pescitelli and Bari recently published a comprehensive review compiling available data on VCD enhancement and classifying the systems that exhibit this phenomenon 22. They combined the systems with an open-shell metal center into a separate group, which contains mononuclear and multinuclear transition metal complexes, lanthanide complexes, and heme proteins. In most cases, the enhancement can be attributed to the coupling of vibrational and electronic excitations; however, other mechanisms can also come into play  22. Sérgio et al. have reported enhanced VCD spectrum for the ligand attached to paramagnetic Co(II)  23, 24, where these systems show enhanced peaks up to two orders of magnitude higher than the usual VCD peaks. They showed that isostructural Co(III) complexes do not exhibit this intensity amplification, providing experimental evidence that the observed enhancement originates from vibronic coupling with low-lying excited states (LLES). In complexes with Schiff base ligands, a sign reversal effect was observed along with the enhanced VCD signals, and this phenomenon was also attributed to the presence of low-lying excited states in a series of studies by Pescitelli 25, 26, 27. Enhancement due to LLES has also been observed in M(III) complexes: for example, Ru(III)(acac)3 studied by More et al. 28, and a tenfold enhancement was reported for [Co(III)stien(biur)2]NBu4 29.
The enhanced intensity provides two practical benefits. First, the overall sensitivity of VCD improves significantly, for example, Lorenzo et al. applied this mechanism to enable a rapid and sensitive chirality recognition in $\alpha$-amino acids 30. Second, the enhancement is distance-dependent; Domingos et al. demonstrated that amplification occurs locally within specific regions of the molecule, providing a way to “zoom in” on particular parts of a biomolecular system 31. However, in order to be able to use these potential applications, a reliable computational approach is needed to analyze and predict the observed enhanced VCD spectra.
Nafie developed the theoretical description of these phenomena, employing a sum-over-states (SOS) formalism that incorporates coupling between vibrational normal modes and LLES 32. Although Nafie’s VCD vibronic theory  32 provides a rigorous framework for vibronic enhancement, it has not been implemented in major quantum chemical software packages, and to date there have been no published ab initio calculations that successfully reproduce enhanced VCD spectra. Tomeček and Bour suggested an alternative approach and studied Co(II)-(-)(sparteine)Cl2 and bis[(S)-N-(1-phenylethyl)salicylaldiminato]$\Delta$-Co(II) complexes. They employed a Herzberg-Teller expansion of the electronic wave function and incorporated nonadiabatic contributions in a model Hamiltonian, followed by full diagonalization of Hamiltonian on the basis of adiabatic wave functions. Although their approach demonstrated clear amplitude amplification, it was still unable to capture the correct signs for several VCD bands. For example, in Co(sp)Cl2, their calculated spectrum reproduced the overall band shape reasonably well, but the predicted enhancement remained roughly three times smaller than the experimental value. In the case of bis[(S)-N-(1phenylethyl)salicylaldiminato]$\Delta$-cobalt(II), their method could not generate the characteristic monosignate VCD profile observed in the experiment.
In this work, we adopt Nafie’s vibronic coupling theory of VCD and argue that previous attempts to predict enhanced VCD failed because existing quantum chemistry methods cannot describe low-lying excited states with the required accuracy. We propose a computational protocol to address these limitations and demonstrate its effectiveness for Me(II)–(--)-sparteine complexes. The article is organized as follows. In Section 2, we review Stephens’ MFP formalism for VCD. Nafie’s vibronic coupling theory is summarized in Section 3. Computational details are provided in Section 4. In Section 5, we present and discuss our results for Co(II) and Ni(II) sparteine complexes. We therein show that, despite the uncertainties in excitation energy prediction by state-of-the-art quantum chemical methods and neglect of Franck-Condon effects, reliable identification of absolute chirality is possible. This opens the way for application of VCD for structure determination of a much broader range of compounds than has been possible with MFP.
2 Magnetic Field Perturbation theory (MFP)
In Stephens MFP formalism for VCD 11, the VCD rotational strength $R_{i}$ is written as 11, 33, 34,
$R_{i}=\textup{Im}[\mathbf{E_{i}}^{\textup{tot}}\cdot{\mathbf{M_{i}}}^{\textup{tot}}],$ (1)
where $i$ is the $i^{th}$ normal mode, $\mathbf{E}^{\textup{tot}}$ and ${\mathbf{M}}^{\textup{tot}}$ are the total electric and magnetic dipole transition moments, respectively. The $\mathbf{E}^{\textup{tot}}$ and ${\mathbf{M}}^{\textup{tot}}$ vectors are defined as follows,
$E^{\textup{tot}}_{\beta,i}={\Bigg{(}{\frac{\hbar}{{\omega}_{i}}}\Bigg{)}}^{\frac{1}{2}}\sum_{\lambda\alpha}P_{\alpha\beta}^{\lambda}S_{\lambda\alpha,i},$ (2)
$M^{\textup{tot}}_{\beta,i}=-{({2{\hbar}^{3}{\omega}_{i}})}^{\frac{1}{2}}\sum_{\lambda\alpha}A_{\alpha\beta}^{\lambda}S_{\lambda\alpha,i},$ (3)
where $\mathbf{S}$ is the transformation matrix from Cartesian ($\lambda\alpha$) to normal ($i$) nuclear coordinates, $\mathbf{P}$ and $\mathbf{A}$ are the atomic polar tensor (APT) and the atomic axial tensor (AAT), respectively. $\omega_{i}$ is the harmonic angular frequency of the $i^{th}$ mode, $\lambda$ is used to enumerate the nuclei and $\alpha$ and $\beta$ indicate Cartesian components of either the nuclear displacement vectors or the electromagnetic field, respectively. The APT and AAT tensors can be written as sum of electronic and nuclear contributions. The APT tensor is then given as,
$P_{\alpha\beta}^{\lambda}=E_{\alpha\beta}^{\lambda}+N_{\alpha\beta}^{\lambda},$ (4)
with the electronic contribution
$E_{\alpha\beta}^{\lambda}={\frac{\partial\langle\Psi_{g}(\mathbf{R})|\mu_{\beta}|\Psi_{g}(\mathbf{R})\rangle}{\partial R_{\lambda\alpha}}\Bigg{|}}_{\mathbf{R}=\mathbf{R}_{0}},$ (5)
and the nuclear contribution
$N_{\alpha\beta}^{\lambda}=eZ_{\lambda}{\delta}_{\alpha\beta}.$ (6)
The AAT tensor is given as follows,
$A_{\alpha\beta}^{\lambda}=I_{\alpha\beta}^{\lambda}+J_{\alpha\beta}^{\lambda},$ (7)
where $\mathbf{I}^{\lambda}$ is the electronic contribution and $\mathbf{J}^{\lambda}$ is the nuclear contribution. The electronic contribution is given as the overlap of the perturbed wave functions with respect to nuclear displacements and with respect to magnetic field,
$I_{\alpha\beta}^{\lambda}={\Bigg{\langle}\frac{\partial\Psi(\mathbf{R},\mathbf{B})}{\partial R_{\lambda\alpha}}\Bigg{|}\frac{\partial\Psi(\mathbf{R},\mathbf{B})}{\partial B_{\beta}}\Bigg{\rangle}\Bigg{|}}_{\mathbf{R}=\mathbf{R}_{0},\mathbf{B}=0},$ (8)
with a nuclear contribution given by
$J_{\alpha\beta}^{\lambda}=i\frac{eZ_{\lambda}}{4\hbar c}\sum_{\gamma}{\epsilon}_{\alpha\beta\gamma}R_{\lambda\gamma}^{0}$ (9)
where ${\epsilon}_{\alpha\beta\gamma}$ is the Levi-Civita symbol.
3 Nafie’s vibronic theory of VCD with LLES
Stephens’ efficient MFP formalism has been implemented in many of the quantum chemistry programs 11, 33, 5 but does not account for the vibronic coupling with LLESs. Equations to express this coupling in terms of additional LLESs contributions to the APT and the AAT tensors have been derived by Nafie 32 and will be employed here. We refer to this paper for the full derivation and will only briefly discuss the SOS formulation of the APT and the AAT tensors.
In Nafie’s SOS formulation of vibronic coupling, the APT tensor ($E_{\alpha\beta}^{A}(\omega_{a})$) is written as
$E_{\alpha\beta}^{\lambda}(\omega_{a})=2\sum_{e\neq g}\Bigg{[}\frac{\omega_{eg}^{2}}{\omega_{eg}^{2}-\omega_{a}^{2}}\Bigg{]}\langle{\Psi_{g}}|\mu_{\beta}|\Psi_{e}\rangle\langle\Psi_{e}|{\frac{\partial\Psi_{g}}{\partial R_{\lambda\alpha}}\rangle\Bigg{|}_{\mathbf{R}=\mathbf{R}_{0}}}.$ (10)
On the RHS of Eq. 10, ${\Psi_{g}}$ is the ground state electronic wave function and ${\Psi_{e}}$ refers to electronically excited-state wave functions. The frequency corresponding to a vertical electronic excitation energy is denoted as $\omega_{eg}$ while the normal mode frequency is given as $\omega_{a}$. In this formalism, the APT is dependent on the normal mode frequency, but it can be related to the one defined by Stephens by writing Eq. 10 in terms of two contributions,
$E_{\alpha\beta}^{\lambda}(\omega_{a})$ $=2\sum_{e\neq g}\Bigg{[}1+\frac{\omega_{a}^{2}}{\omega_{eg}^{2}-\omega_{a}^{2}}\Bigg{]}\langle{\Psi_{g}}|\mu_{\beta}|\Psi_{e}\rangle\langle\Psi_{e}|{\frac{\partial\Psi_{g}}{\partial R_{\lambda\alpha}}\rangle\Bigg{|}_{\mathbf{R}=\mathbf{R}_{0}}}$ (11)
$=2\sum_{e\neq g}\langle{\Psi_{g}}|\mu_{\beta}|\Psi_{e}\rangle\langle\Psi_{e}|{\frac{\partial\Psi_{g}}{\partial R_{\lambda\alpha}}\rangle\Bigg{|}_{\mathbf{R}=\mathbf{R}_{0}}}$
$+{\underbrace{2\sum_{e\neq g}\frac{\omega_{a}^{2}}{\omega_{eg}^{2}-\omega_{a}^{2}}\langle{\Psi_{g}}|\mu_{\beta}|\Psi_{e}\rangle\langle\Psi_{e}|{\frac{\partial\Psi_{g}}{\partial R_{\lambda\alpha}}\rangle\Bigg{|}_{\mathbf{R}=\mathbf{R}_{0}}}}_{\textup{enhancement term}}}.$
The first term in the RHS of Eq. 11, is equivalent to the APT defined in the MFP formalism,
$2\sum_{e\neq g}\langle{\Psi_{g}}|\mu_{\beta}|\Psi_{e}\rangle\langle\Psi_{e}|{\frac{\partial\Psi_{g}}{\partial R_{\lambda\alpha}}\rangle\Bigg{|}_{\mathbf{R}=\mathbf{R}_{0}}}$ (12)
$=2\langle{\Psi_{g}}|\mu_{\beta}|{\frac{\partial\Psi_{g}}{\partial R_{A\alpha}}\rangle\Bigg{|}_{\mathbf{R}=\mathbf{R}_{0}}}-2\langle{\Psi_{g}}|\mu_{\beta}|\Psi_{g}\rangle\langle\Psi_{g}|{\frac{\partial\Psi_{g}}{\partial R_{\lambda\alpha}}\rangle\Bigg{|}_{\mathbf{R}=\mathbf{R}_{0}}}$
$={\frac{\partial\langle\Psi_{g}|\mu_{\beta}|\Psi_{g}\rangle}{\partial R_{\lambda\alpha}}\Bigg{|}}_{\mathbf{R}=\mathbf{R}_{0}},$
since $\langle\Psi_{g}|\frac{\partial\Psi_{g}}{\partial R_{\lambda\alpha}}\rangle=0$ and the electron dipole moment operator carries no direct dependence on nuclear coordinates. The enhancement can thus be solely attributed to the second term of Eq. 11 which depends both on the normal mode frequency and the electronic excitation energy.
Like the APT, the AAT in the SOS formalism depends on both the normal mode frequency and the electronic excitation energy. It takes the form
$I_{\alpha\beta}^{\lambda}(\omega_{a})$ $=\sum_{e\neq g}\Bigg{[}\frac{\omega_{eg}^{2}}{\omega_{eg}^{2}-\omega_{a}^{2}}\Bigg{]}\frac{\langle\frac{\partial\Psi_{g}}{\partial R_{\lambda\alpha}}|\Psi_{e}\rangle\langle\Psi_{e}|m_{\beta}|\Psi_{g}\rangle}{{\omega}_{eg}}\Bigg{|}_{\mathbf{R}=\mathbf{R}_{0}},$ (13)
where $\mathbf{m}$ is the magnetic dipole operator and we have multiplied the original definition of Nafie by a factor of $\frac{i}{2}$ to be consistent with the one of Stephens8. Eq. 13 can be partitioned as,
$I_{\alpha\beta}^{A}(\omega_{a})$ $=\sum_{e\neq g}\Bigg{[}1+\frac{\omega_{a}^{2}}{\omega_{eg}^{2}-\omega_{a}^{2}}\Bigg{]}\frac{\langle\frac{\partial\Psi_{g}}{\partial R_{A\alpha}}|\Psi_{e}\rangle\langle\Psi_{e}|m_{\beta}|\Psi_{g}\rangle}{{\omega}_{eg}}\Bigg{|}_{\mathbf{R}=\mathbf{R}_{0}}$ (14)
$=\sum_{e\neq g}\frac{\langle\frac{\partial\Psi_{g}}{\partial R_{A\alpha}}|\Psi_{e}\rangle\langle\Psi_{e}|m_{\beta}|\Psi_{g}\rangle}{{\omega}_{eg}}\Bigg{|}_{\mathbf{R}=\mathbf{R}_{0}}$
$+{\underbrace{\sum_{e\neq g}\Bigg{[}\frac{\omega_{a}^{2}}{\omega_{eg}^{2}-\omega_{a}^{2}}\Bigg{]}\frac{\langle\frac{\partial\Psi_{g}}{\partial R_{A\alpha}}|\Psi_{e}\rangle\langle\Psi_{e}|m_{\beta}|\Psi_{g}\rangle}{{\omega}_{eg}}\Bigg{|}_{\mathbf{R}=\mathbf{R}_{0}}}_{\textup{enhancement term}}}.$
The first term of Eq. 14 can be related to Eq. 8:
$I_{\alpha\beta}^{A}$ $={\Bigg{\langle}\frac{\partial\Psi_{g}}{\partial R_{A\alpha}}\Bigg{|}\frac{\partial\Psi_{g}}{\partial B_{\beta}}\Bigg{\rangle}\Bigg{|}}_{\mathbf{R}=\mathbf{R}_{0},\mathbf{B}=0}$ (15)
$=\sum_{e\neq g}\frac{\langle{\frac{\partial\Psi_{g}}{\partial R_{A\alpha}}|\Psi_{e}\rangle\langle\Psi_{e}|m_{\beta}|\Psi_{g}\rangle}}{{\omega}_{eg}}\Bigg{|}_{\mathbf{R}=\mathbf{R}_{0}},$
and the second term of Eq. 14 represents the enhancement.
These equivalences make it possible to compute the VCD rotational strength as a sum of standard and enhancement contributions and thereby allows for detailed analysis. We define
$R_{i}^{\textup{enh}}=\textup{Im}[\mathbf{E}_{i}^{\textup{tot}}\cdot\mathbf{M}_{i}^{\textup{enh}}+\mathbf{E}_{i}^{\textup{enh}}\cdot\mathbf{M}_{i}^{\textup{tot}}]$ (16)
$E_{i,\beta}^{\textup{enh}}=2{\Bigg{(}{\frac{\hbar}{{\omega}_{i}}}\Bigg{)}}^{\frac{1}{2}}\sum_{e\neq g}\frac{\omega_{i}^{2}}{\omega_{eg}^{2}-\omega_{i}^{2}}\Bigg{[}\langle{\Psi_{g}}|\mu_{\beta}|\Psi_{e}\rangle\langle\Psi_{e}|\frac{\partial\Psi_{g}}{\partial R_{i}}\rangle\Bigg{]}_{\mathbf{R}=\mathbf{R}_{0}}$
$M_{i,\beta}^{\textup{enh}}=-{({2{\hbar}^{3}{\omega}_{i}})}^{\frac{1}{2}}\sum_{e\neq g}\frac{\omega_{i}^{2}}{\omega_{eg}^{2}-\omega_{i}^{2}}\Bigg{[}\frac{\langle{\Psi_{g}}|m_{\beta}|\Psi_{e}\rangle}{\omega_{eg}}\langle\Psi_{e}|{\frac{\partial\Psi_{g}}{\partial R_{i}}\rangle\Bigg{]}_{\mathbf{R}=\mathbf{R}_{0}}}$
The enhancement corrections can be calculated separately and require evaluation of excitation energies, electric-dipole transition, magnetic dipole transition for the LLES as are readily available in most quantum chemistry codes. The required non-adiabatic coupling vectors (NACVs) $\langle\Psi_{e}|\frac{\partial\Psi_{g}}{\partial R_{A\alpha}}\rangle$ can also be obtained from a number of programs.
Nafie’s formulation employs the vertical approximation, in which the ground and excited state potential energy surfaces are assumed to have the same shape and differ only by a constant energy shift. Within this approximation, one assumes $\phi^{a}_{ev}=\phi^{a}_{gv}$ for all vibrational states ($v$) of elecronic ground ($g$) and excited states ($e$). Consequently, overlap integrals of the form $\braket{\phi^{a}_{g0}}{\phi^{a}_{ev}}$ are zero for all transitions except the fundamental $0\rightarrow 0$ transition. As we show in the results section, however, this vertical approximation is not holding for all low-lying excited states of sparteine complexes and can be one of the reasons of deviations of the calculated and observed spectra.
4 Computational details
We study the low-lying excited states effects on the VCD spectra of the open-shell complexes Co(II)(sp)Cl2 and Ni(II)(sp)Cl2 . The cobalt (d7 electronic configuration) and nickel (d8 electronic configuration) atoms form high-spin open-shell, four-coordinate complexes with the bidentate chiral (-)-sparteine ligand with the absolute configuration of 6R, 7S, 9S, 11S ( Figure 1). They can be compared with analogous zinc complexes where the electronic structure is that of a closed shell configuration with no unpaired spins, significantly higher excitation energies, and therefore negligible enhancement effects. To gain more insight about the electronic structure of the excited state, we also study model complexes with $C_{2v}$ symmetry. These model complexes are constructed by truncating and symmetrizing the optimized sparteine complexes while preserving bond distances around the transition metal.
Figure 1: Co(sp)Cl2 structure optimized on BP86-D3BJ/def2-TZVP level with CPCM treatment of CHCl3 solvent effects (left) and the $C_{2v}$ Co(NH3)2Cl2 model structure.
According to previous studies  35, the sparteine complexes are inherently rigid as a result of the ligand structure itself. To provide an independent verification of this finding, we use CREST  36 within the Amsterdam Modeling Suite (AMS)  37 to generate initial conformers of Zn(sp)Cl2 employing the GFN1-xTB method  38. Conformers within 5 kcal/mol of the lowest energy structure are subsequently reoptimized at the DFT level using BP86 39, 40-D3BJ 41, 42/def2-TZVP with the conductor-like polarizable continuum model (CPCM)  43 to model CHCl3 solvent effects. All the IR and MFP VCD are performed on the same level of theory for closed-shell Zn(sp)Cl2 and open-shell Co(sp)Cl2 and Ni(sp)Cl2 complexes.
We calculate excited states of open-shell complexes using CASSCF and TDDFT, including dipole transition moments (EDTM/MDTM) and non-adiabatic couplings. TDDFT calculations are performed with (ORCA 6.0) 44. We use def2-TZVP 45 basis set with def2/JK 46 auxiliary sets and extract electric and magnetic transition dipole moments (EDTM/MDTM) and non-adiabatic couplings for further enhancement calculations. CASSCF calculations are performed with both ORCA 6.0 and Openmolcas (v23.06) 47. In ORCA, we perform state-averaged (SA-) and state-specific (SS-) CASSCF on sparteine and model $C_{2v}$ complexes with the def2-TZVP basis set and the def2 / JK auxiliary fitting. OpenMolcas is employed specifically for its implementation of non-adiabatic coupling vectors at the SA-CASSCF level; there, we use the ANO-RCC-VTZP basis on the metal  48, N, and Cl atoms  49, and the minimal ANO-RCC-MB basis on C and H.
We calculate enhanced VCD spectra using our developed Python implementation of Nafie’s vibronic coupling theory described in Section 3. To quantify the agreement of the calculated IR and VCD spectra with the observed ones we use Shen  50 simIR and simVCD metrics. To this end, we digitized the experimental IR and VCD spectra for all sparteine complexes from the thesis of He 51 using a Python script. A frequency scaling factor is obtained by maximizing the simIR between the computed and experimental IR spectra over the 950–1500 cm-1 window, and then the same factor is applied to the calculated VCD spectra.
5 Results and discussion
5.1 Magnetic field perturbation calculations
Zn(II) has a closed-shell $d^{10}$ configuration with high-lying excited states, so enhancement effects for Zn(sp)Cl2 are negligible and the MFP approximation is fully justified. As a result, this complex serves as a good benchmark for MFP-based VCD calculations. The IR spectra of all three studied M(sp)Cl2 (M = Zn, Co, Ni) are almost identical, so the Zn(sp)Cl2 complex can be used to select the most appropriate DFT functional and to evaluate conformational effects. Conformational analysis reveals an energy difference between the two lowest energy conformers of Zn(sp)Cl2 of 4.2 kcal/mol which exceeds the value 3 kcal/mol typically used to select contributing conformers. This is in agreement with previous studies of the sparteine ligand  35 who also found Me(II)(sp)Cl2 complexes to exist predominantly in a single conformation and makes the systems ideal for analysing their VCD spectra in detail.
Several benchmark studies have shown that the generalized gradient approximation (GGA) functional BP86 39, 52 outperforms B3LYP 53 in reproducing experimental equilibrium geometries of small first-row transition metal complexes 54, 55. Minenkov et al.56 furthermore reported that B3LYP systematically overestimates the metal-ligand bonds in transition metal complexes of the first and second row. A more recent benchmark for metal–dinitrogen complexes 57 further showed that DFT performance is highly sensitive to the nature of the N2 coordination mode at the metal center. We compare the performance of three DFT functionals (BP86, B3LYP, PBE0 58 with and without Grimme’s D3(BJ) dispersion correction  41, 42 in the gas phase and with CPCM of CHCl3  43 (see detailed plots in the Supporting Information).
Figure 2 shows IR and MFP-VCD of Zn(sp)Cl2 with the considered functionals. The addition of the dispersion correction improves both simIR and simVCD for BP86 and B3LYP (see Figure S1 in SI). BP86-D3BJ/def2-TZVP with CPCM demonstrates the best agreement with the experiment with good similarities for both IR and VCD (simIR = 0.88, simVCD = 0.44). Typically, simVCD values larger than 0.4 are considered to be secure to assign the absolute configuration 59. The calculated IR spectra of open-shell Co(sp)Cl2 and Ni(sp)Cl2Figure 3) at the same level of theory have a high similarity to the experiment. This confirms the suitability of BP86-D3BJ/def2-TZVP with CPCM to optimize the ground state structure and compute IR spectra for these compunds.
Figure 2: Comparison of VCD (top) and IR (bottom) spectra of Zn(sp)Cl2. Experimental spectra are given in black, the spectra are calculated with CPCM of CHCl3 with B3LYP/def2-TZVP (red, simVCD = 0.16), PBE0/def2-TZVP (green, simVCD = 0.28), BP86/def2-TZVP (blue, simVCD = 0.44).
Figure 3: Co(sp)Cl2 calculated (blue, simIR = 0.88) and observed (black) IR spectra (left). Ni(sp)Cl2 calculated (red, simIR = 0.85) and observed (black) IR spectra (right). All calculations are performed with BP86-D3BJ/def2-TZVP with CPCM of CHCl3.
5.2 Excited state calculations
We next investigate the excitation energies and the nature of LLES for the Co(II)(sp)Cl2 and Ni(II)(sp)Cl2 complexes with TDDFT and SA-CASSCF. The contributions of each excited state to the enhanced VCD spectra depend on the EDTM, MDTM, excitation energy, nonadiabatic coupling matrix elements, and vibrational frequencies, so the quality of all these quantities is important to obtain a reliable estimate of the enhancement that can be expected in VCD.
According to experimental studies, pseudotetrahedral M(sp)Cl2 complexes exhibit low-lying, dipole-forbidden $d-d$ transitions that underlie their enhancement effect 21. Our SA-CASSCF calculations on Co(II)(sp)Cl2 and Ni(II)(sp)Cl2 confirm the dominant $d$ character of active orbitals with Löwdin orbital compositions 94% and higher (see Figure 4). In the main text, we present results for minimal active spaces (7,5) for Co(II) and (8,5) for Ni(II), while data for the larger active spaces for Co(sp)Cl2 is provided in the Supporting Information (Table S2). Since the transitions of interest are highly localized, we also construct a smaller model complex, M(II)(NH3)2Cl2, with $C_{2v}$ symmetry. This higher symmetry allows state specific CASSCF calculations, thereby avoiding possible artifacts introduced by orbital averaging in SA-CASSCF.
Figure 4: SA-CASSCF(7, 5) natural active orbitals of Co(sp)Cl2 with occupation numbers (n). Löwdin orbital compositions are shown for atomic orbitals contributing more than 5%.
Co(sp)Cl2 SA-CASSCF(7, 5) calculations show that the first three excited states lie approximately 0.2 eV below the next higher state (see  Figure 5). Since each state’s contribution scales inversely with $w_{eg}-w_{a}$, only those excited states whose energies nearly coincide with a vibrational mode yield significant enhancement. Consequently, contributions from the fourth and higher excited states are effectively negligible. For Ni(sp)Cl2, there is already a significant gap between the first and second excited states (see  Figure 5) and for this complex only the first excited state needs to be considered in the enhancement calculations. In  Figure 5, we compare the excitation energies computed with representative GGA, hybrid, and hybrid meta-GGA functionals. Pure GGAs substantially overestimate the energies of the low-lying states for a high-spin complex. Introducing Hartree–Fock exchange systematically improves the agreement: B3LYP (20% HF exchange) reduces the error, and BHandH (50% HF exchange) performs even better. In contrast, the hybrid meta-GGA M06 60 yields a too low first excitation energy relative to the SA-CASSCF benchmark values.
Figure 5: Excitation energies for the first 6 excited states calculated with TDDFT and SA-CASSCF in Orca 44 for Co(sp)Cl2 (left) and Ni(sp)Cl2 (right). We use the symmetry labels from C2v model (see Table 1 complex to distinguish the excited states for non-symmetric sparteine. The functionals are ordered from GGA (BP86) to hybrid functionals with increasing Hartree-Fock exchange: B3LYP (20%), PBE0 (20%), BHandH (50%). The Hybrid Meta-GGA functional M06 yields a too low first excitation energy, and is put to the right of the reference results to ease the comparison.
Table 1 summarizes CASSCF/NEVPT2 excitation energies for the sparteine and symmetric model complexes. Notably, the SA-CASSCF(7,5) results for Co(sp)Cl2 and Co(NH3)2Cl2 are nearly identical, confirming the highly localized character of these transitions. For the $C_{2v}$ complex the ground (${}^{4}A_{2}$) and excited states (${}^{4}A_{1}$, ${}^{4}B_{1}$ and ${}^{4}B_{2}$) each belong to different irreducible representations, which permits state specific CASSCF optimizations. Here, the first two excited states are almost degenerate, while the third lies slightly higher in energy. For Ni(II)(NH3)3Cl2 complex, the ground state is ${}^{3}B_{2}$, and its three lowest excited states are ${}^{3}A_{2}$, ${}^{3}B_{1}$, and ${}^{3}B_{2}$. Overall, the close agreement between the SS-CASSCF energies of the high-symmetry model and the SA-CASSCF results confirms that the state-averaged orbitals provide a reliable description of these local excitations.
Table 1: Vertical (v) and adiabatic (a) excitation energies of sparteine and model complexes calculated with CASSCF method. NEVPT2 corrected energies are given in brackets. All excitation energies are given in eV.
Co(II)(NH3)2Cl2 Co(II)(sp)Cl2
Symmetry SS-CASSCF (v) (NEVPT2) SA-CASSCF (v) (NEVPT2) SA-CASSCF (v) (NEVPT2) SA-CASSCF (a) (NEVPT2)
${}^{4}A_{2}$rightarrow ${}^{4}A_{1}$ 0.35 (0.39) 0.34 (0.44) 0.34 (0.42) 0.22 (0.34)
${}^{4}A_{2}$rightarrow ${}^{4}B_{1}$ 0.34 (0.42) 0.37 (0.45) 0.37 (0.48) 0.32 (0.51)
${}^{4}A_{2}$rightarrow ${}^{4}B_{2}$ 0.41 (0.50) 0.41 (0.52) 0.42 (0.53) 0.26 (0.45)
Ni(II)(NH3)2Cl2 Ni(II)(sp)Cl2
Symmetry SS-CASSCF (v) (NEVPT2) SA-CASSCF (v) (NEVPT2) SA-CASSCF (v) (NEVPT2) SA-CASSCF (a) (NEVPT2)
${}^{3}B_{2}$rightarrow ${}^{3}A_{2}$ 0.30 (0.35) 0.29 (0.37) 0.30 (0.39) 0.25 (0.34)
${}^{3}B_{2}$rightarrow ${}^{3}B_{1}$ 0.48 (0.52) 0.43 (0.57) 0.47 (0.61) 0.45 (0.72)
${}^{3}B_{2}$rightarrow ${}^{3}A_{1}$ 0.63 (0.72) 0.62 (0.83) 0.70 (0.92) 0.53 (0.91)
One key approximation in the Nafie framework that we already mentioned when discussing his theory is the vertical approximation. To check how this approximation can affect results, we show in the last column of Table 1 the adiabatic excitation energies of the sparteine complexes. For the ground state we retain the BP86 geometry, as it yields geometries in best agreement with experimental data (see Figure S2 in SI). For the excited states, however, we decided to optimize geometries using the BHandH functional which provides a state description more closely resembling that obtained from SA-CASSCF than is obtained with TDDFT employing BP86 (see additional results in Table S3 in SI). We thereby assume that, since the calculation of the TDDFT excited-state gradients requires evaluation of many terms shared with a non-adiabatic couplings (NACs) calculation within the linear response framework, a functional that yields more accurate NACs is likely to provide a more reliable description of excited-state gradients.
In Co(sp)Cl2, excited state relaxation lowers the excitation energy by more than 0.10 eV, which is significant in case of such small energies. This shows the limitation of using a vertical approximation, and indicates that inclusion of Franck-Condon factors is likely necessary to achieve quantitatively accurate enhanced VCD predictions by fully ab initio calculation. As this will lead to impractically large computational demands, we will in the next sections demonstrate that a semi-empirical approach can be used instead.
5.3 SA-CASSCF calculations of the enhancement
We compute the enhanced VCD spectra using our Python implementation of Equation 16. The APT and AAT tensors are evaluated at the BP86-D3BJ/CPCM level, and corrections to these tensors are obtained from SA-CASSCF calculations in OpenMolcas averaged over 4 states. The results of VCD calculations for Co(sp)Cl2 are given in Figure 6. Incorporation of three excited states with unmodified vertical excitation energies produces a slight increase in simVCD compared to the MFP, but fails to reproduce the pronounced amplification observed experimentally. Given the uncertainty due to the accuracy of the computed excitation energy and the effect of the vertical approximation, we decided therefore to treat excited state energies as tunable parameters which are optimized to maximize the simVCD value. This is inspired by a similar approach used by some of us in earlier work to deal with uncertainties in determining Boltzmann weights in conformational averaging59. In our current approach the transition dipole moments and non-adiabatic couplings are all unchanged and taken directly from the SA-CASSCF calculations, having only the three excited state energies as tunable parameters. Using this approach we can achieve reliable agreement with the experimental spectra with a simVCD value of 0.69. The energies of the first two excited states converge to an identical value of 0.23 eV, which is approximately 0.1 eV lower than the SA-CASSCF value. This deviation is well within the expected error of this method. The energy of the third state did not change by optimization and remained at 0.42 eV, thereby reducing the number of required tunable parameters required to just two.
Figure 6: Enhanced VCD spectra of Co(sp)Cl2 calculated using MFP (red, simVCD = 0.15), SA-CASSCF (green, simVCD = 0.30), and optimized excited state energies (blue, simVCD = 0.69) are demonstrated on the left. Dependence of the VCD intensity on the energy value of the first two excited states is shown on the right.
Figure 6 illustrates the sensitivity of the simulated VCD enhancement to the energies of the first two excited states. The calculated peaks are slightly shifted relative to the experimental spectrum because we applied only a single frequency scaling factor from the IR calculations. Even 0.02 eV changes in excitation energies significantly affects the amplification of the 1300–1500 cm-1 bands. Such small energy difference far exceeds the typical uncertainty of even high-level quantum chemistry methods. Consequently, fitting low-lying excitation energies to experimental VCD data would be practical route to reproduce the observed enhancement.
Figure 7: Individual contributions of the excited states to the enhancement for Co(sp)Cl2 with the optimized energies. The first two states have both excitation energies 0.23 eV and give similarity 0.38 (dark blue) and 0.29 (blue) respectively. The excitation energy of the third state is 0.42 eV and it’s contribution almost zero (light blue). The grey line corresponds to the excitation energy of the third state shifted to 0.23 eV, the simVCD is -0.11 eV in this case.
The individual contributions of each excited state to the enhanced VCD spectrum are shown in  Figure 7. Considering only the first or only the second excited states increases simVCD values by 0.38 and 0.29, respectively. In contrast, employing the third excited state (at 0.42 eV) gives virtually no similarity to the observed spectrum (simVCD = –0.01, light-blue trace), indicating that its energy lies above the resonance window required for enhancement. When the third state’s energy is artificially shifted downward to 0.23 eV (gray trace), its simVCD even has an inverted sign, giving further confirmation that it does not contribute to the experimentally observed spectrum. From these results, we conclude that only the two lowest excited states drive the vibronic enhancement in Co(sp)Cl2.
For Ni(sp)Cl2 the enhanced spectra with unmodified SA-CASSCF energies also show only slight enhancement compared to MFP spectra  Figure 8). Optimization of the excitation energies including 3 lowest excited states gives a simVCD value of 0.51 with tuned energies of 0.24, 0.31, and 0.37 eV. Including only the lowest state in the optimization yields virtually identical spectra, indicating that in this case only one state contributes to the enhancement. This observation is consistent with the calculated SA-CASSCF excitation energies which predict the second and third excited states to lie significantly higher than the first. After the optimization the calculated Ni(sp)Cl2 spectrum shows one important difference with the experimentally observed one: in the calculated enhanced spectra an intense negative peak appears around 1300 cm-1. This discrepancy already appears in the MFP spectra and may for instance arise from artifacts due to the implicit modeling of solvent effects. If we calculate the spectra similarity excluding this 1300-1320 cm-1 region, where the discrepancy is observed the simVCD is increased to 0.57.
Figure 8: Enhanced VCD spectra of Ni(sp)Cl2 calculated with MFP (red, simVCD = 0.04), SA-CASSCF (green, simVCD = 0.22) and SA-CASSCF with optimized energies including one (blue, simVCD = 0.51) and three states (dark blue, simVCD = 0.51).
We conclude that SA-CASSCF, with tuning of the lowest excitation energies, can yield very reasonable enhanced VCD spectra for the sparteine complexes. However, SA-CASSCF calculations of non-adiabatic couplings will present a serious bottleneck for applying this approach to larger systems. To be able to overcome this computational bottleneck, we will next explore TDDFT as a more scalable alternative to model VCD intensity enhancement.
5.4 TDDFT calculations of the enhancement
Most density functional approximations that we studied predict excitation energies that are too high to contribute to the enhancement (see additional results in the Table S4 and S5 in SI). To remedy this method artifact, we optimize the excitation energies in the same way as in the SA-CASSCF approach, while retaining the transition dipole moments and non-adiabatic couplings from the TDDFT calculations. BHandH and B3LYP yield nearly identical simVCD values; here, we report the BHandH results since they give the largest simVCD (see  Figure 9). Data for B3LYP and other DFAs is available in the Supporting Information.
Figure 9: Enhanced VCD spectra of Co(sp)Cl2 calculated with BHandH functional including two excited states (red, simVCD = 0.51) and three excited states (blue, simVC = 0.67) on the left. Individual contributions of the first three excited states to the enhancement for Co(sp)Cl2 from BHandH calculation with similarities 0.20, 0.25 and 0.16.
For the Co(sp)Cl2 complex, optimization of three excited states energies produces close values around 0.23 eV. This differs from the SA-CASSCF results, where the third state remained higher and did not contribute to the enhancement. We link such differences to the non-adiabatic couplings (NACs) (see Figure 10). The only significant coupling values correspond to atoms directly bonded to the transition metal, again confirming the distance-dependence of the enhancement effect. The NACs between the ground and second excited states are very close for BHandH and SA-CASSCF. More significant deviations are observed for the first and the third state. In the third state, the couplings involving the N atoms are noticeably stronger with BHandH, which likely enables it’s contribution to the enhancement. We find that the non-adiabatic coupling vectors computed with B3LYP closely match those obtained with the BHandH DFA, which accounts for the nearly identical enhanced VCD profiles predicted by both methods. NACs, difference plots and mean errors for other DFAs can be found in the SI.
(a) NACs between the ground and first excited state
(b) NACs between the ground and second excited state
(c) NACs between the ground and third excited state
Figure 10: Non-adiabatic couplings between ground and first three excited states of Co(sp)Cl2 calculated with BHandH and SA-CASSCF for atoms around Co(II).
When we restrict the optimization to just the two lowest excited states, the agreement with experiment remains strong: the overall simVCD falls to 0.54 (versus 0.68 when all three states are adjusted), yet the spectral profile is still well reproduced. Most of the discrepancy arises in the 1250–1400 cm-1 region, and in a more pronounced negative peak around 1100 cm-1. The right-hand panel of  Figure 9 demonstrate contributions of each excited state, revealing that the same VCD band can flip sign depending on the state. For instance, the first excited state produces a negative feature near 1300 cm-1, while both the second and third excited states give positive peaks at that frequency. Moreover, different states dominate distinct spectral regions: in the 1500–1200 cm-1 region a superposition of the first and second excited state contributions reproduces the overall band shape, the 1200-1100 cm-1 region is well described by the second state, and finally 1100-1000 cm-1 is closer to the shape of the third state contribution.
For Ni(sp)Cl2, the TDDFT results using the BHandH functional are essentially identical to those obtained with SA-CASSCF: only the first excited state contributes to the enhancement, giving a final simVCD of 0.45 (see  Figure 11) with optimized energy 0.23 eV. The non-adiabatic couplings computed with BHandH closely match those from SA-CASSCF for Ni(sp)Cl2, which explains the close result ( Figure 12). We again observe a large negative peak at 1300 cm-1, similar to SA-CASSCF case.
Figure 11: Enhanced VCD spectra of Ni(sp)Cl2 calculated with BHandH functional with one (red, simVCD = 0.45) and three (blue, simVCD = 0.46) excited states.
(a) NACs between the ground and first excited state
(b) NACs between the ground and second excited state
(c) NACs between the ground and third excited state
Figure 12: Non-adiabatic couplings between ground and first three excited states of Ni(sp)Cl2 calculated with BHandH and SA-CASSCF.
5.5 Assignment of absolute configuration
An critical test of our semi-empirical approach is its ability to assign absolute configuration reliably. To probe this, we inverted the experimental VCD spectrum and repeated the energy optimization procedure. For Co(sp)Cl2 the optimization drives the first excited-state energy down to an unphysically low 0.08 eV, where perturbation theory breaks down (see Figure 13). The other two states converge to 0.37 eV and contribute negligibly. The resulting VCD intensities are dramatically overestimated, especially in the 1100–1000 cm-1 region, which is a clear evidence of a breakdown in the fitted enantiomer spectrum. For Ni(sp)Cl2 the optimized simVCD remains moderately high (0.32) and the spectral shape is closer to the inverted spectrum than in the Co case. However, state analysis reveals that only the third excited state contributes constructively to the fit. Moreover, optimization drives the energy of this state to a low energy of 0.18 eV, while the first two states end up with degenerate energies of 0.37 eV. In general, we show that fitting the spectrum of another enantiomer with the proposed optimization procedure leads to results, which can be readily identified as being erroneous. This makes it possible to rely on the tuned enhanced calculation result for robust and unambiguous assignment of absolute configuration, provided that the set of low-lying states used in the tuning is carefully selected and validated.
Figure 13: Enhanced VCD spectra for the enantiomer of Co(sp)Cl2 (left). The calculated spectrum is obtained using SA-CASSCF with three excited states and optimized excitation energies (red); the enantiomeric VCD spectrum is obtained by inversion (black). Individual contributions of each state are given on the right. The second and the third state have negligible contributions with simVCD -0.05 (red) and 0.02 (purple), respectively.
Figure 14: Enhanced VCD spectra for the enantiomer of Ni(sp)Cl2 (left). The calculated spectrum is obtained using SA-CASSCF states and optimized excitation energies (red); the enantiomeric VCD spectrum is obtain by inversion (black). Individual contributions of each state are given on the right. The first state has a negative similarity -0.12 (dark red); the overall positive similarity is coming from the third state (purple, simVCD = 0.34)
5.6 Understanding the mechanism of enhancement
Another advantage of the formalism presented in this work is the possibility to understand the mechanism of the enhancement in full detail. To this end we note that we can distinguish enhancement due to the magnetic and electric dipole components of the electronic transition moments:
$R_{i}^{\textup{enh}}=\textup{Im}[\mathbf{E}_{i}^{\textup{tot}}\cdot\mathbf{M}_{i}^{\textup{enh}}+\mathbf{E}_{i}^{\textup{enh}}\cdot\mathbf{M}_{i}^{\textup{tot}}]$
We observe for all the calculated enhanced spectra that the electric dipole contribution is negligible. This arises from the small electric dipole transition moments associated with pure $d-d$ transitions. The decomposition of the enhanced spectra into contributions for Co(sp)Cl2 and Ni(sp)Cl2 is demonstrated in Figure 15. We clearly observe the magnetic dipole enhancement dominating the signal, making the accuracy of the APT tensor (from the MFP calculation) and the magnetic dipole transition moments between electronic states more critical than the AAT tensor and the electric dipole transition moments.
Figure 15: Decomposition of enhanced VCD spectra into MFP (red), electric dipole enhancement (green) and magnetic dipole enhancement (blue) for Co(sp)Cl2 (left) and Ni(sp)Cl2 (right) calculated with SA-CASSCF excited states properties and optimized geometries.
6 Conclusions
In this work, we have applied Nafie’s sum-over-states approach to calculate the VCD enhancement in transition metal complexes. We studied rigid sparteine complexes as a proof-of-concept application of vibronic coupling theory. The next step will be to extend this approach to flexible complexes, where conformational variability introduces additional complications.
We find that the enhancement is extremely sensitive to the excitation energies, making it challenging for quantum chemistry methods to predict these values with sufficient accuracy. Rather than pushing for even more accurate excited state calculations, we propose treating these key excitation energies as adjustable parameters. By optimizing just a few values, we can reproduce experimental VCD spectra with high similarity, achieving simVCD over 0.4 for both SA-CASSCF and TDDFT.
At the CASSCF level, only two excited states for the cobalt complex and one for the nickel complex are needed to reproduce the main features of the enhanced VCD spectra, demonstrating that the sum-over-states expansion converges rapidly in these systems. Including different excited states can change the sign of a given vibrational transition, which highlights the importance of selecting the appropriate states in the enhancement calculation.
Our results suggest that TDDFT can perform comparably to CASSCF, provided that an appropriate functional is chosen. In our study, BHandH delivers the best VCD enhancements, likely due to its high fraction of Hartree–Fock exchange, which is crucial for accurately describing $d-d$ transitions in transition metal complexes. In fact, for Ni(sp)Cl2, the non-adiabatic couplings calculated with BHandH closely match those of SA-CASSCF, leading to nearly identical spectra with SA-CASSCF states. For Co(sp)Cl2, we observe some discrepancies in the couplings to the first and third excited states between the two methods, accounting for the slight differences in their respective contributions.
An important observation, noted in earlier studies, is the local character of the VCD enhancement. Here, we show that significant nonadiabatic couplings occur only between the transition metal and its close neighbors: two nitrogens and two chlorines in the sparteine case. Consequently, we assume that for larger systems, one could compute these couplings on a truncated model without substantial loss of accuracy. We plan to explore this idea in future work.
Our findings furthermore demonstrate that problems to computationally reproduce enhanced VCD spectra were due to inaccuracies in just a few key parameters. Making these parameters tunable leads to an efficient and practical scheme to determine absolute configurations by means of enhanced VCD spectra in transition metal systems.
7 Acknowledgements
We acknowledge financial support from NWO in the framework of the Fund New Chemical Innovations (NWO Project Nr.731.014.209) and the National Grow Fund Quantum Delta NL (NWO Project Nr. NGF.1582.22.036). SAJ acknowledges support of the Chemistry & Chemical Engineering Research Center of Iran (CCERCI) in the framework of a student exchange program. We furthermore acknowledge NWO for providing access to Snellius, hosted by SURF through the Computing Time on National Computer Facilities call for proposals.